OpenStudy (anonymous):

Find the x- and y-intercepts and graph the equation plotting the intercepts. -6x - 3y = 27

mathslover (mathslover):

first put x = 0 in the equation to find y intercept : -6(0)-3y=27 solve for y, what do you get?

OpenStudy (anonymous):

x=9? This is so confusing

mathslover (mathslover):

what is -6 * 0 ?

OpenStudy (anonymous):

0

mathslover (mathslover):

right now: -6*0 -3y = 27 -3y =27 right?

OpenStudy (anonymous):

yes

mathslover (mathslover):

now can u solve for y?

mathslover (mathslover):

@kartiksriramk please let me explain the asker... You have already given a direct soln that is against Code of Conduct..

OpenStudy (anonymous):

Oh sorry man.......I'll remove it.

OpenStudy (anonymous):

I'm sorry, I just don't get it.

mathslover (mathslover):

-3y = 27 divide both sides by -3 and then try to solve

OpenStudy (anonymous):

y=-3

OpenStudy (anonymous):

no, y = -9

mathslover (mathslover):

y = - 9 correct

mathslover (mathslover):

now put y = 0 that is solve for x: -6x+3(0) =27

OpenStudy (anonymous):

is it not supposed to be -6x - 3(0)=27?

mathslover (mathslover):

oh yes sorry - is there so solve for x : -6x = 27

OpenStudy (anonymous):

-6x - 3(0) = 27-6x=27?

mathslover (mathslover):

yes.. since -3*0 = 0 hence -6x-3*0=27 -6x=27

OpenStudy (anonymous):

ok so do i divided both sides by -6x?

mathslover (mathslover):

no no divide both sides by -6

OpenStudy (anonymous):

this part is confusing, what am i missing?

mathslover (mathslover):

\[\large{\frac{-6x}{-6}=\frac{27}{-6}}\] \[\large{\frac{\cancel{-6}^1x}{\cancel{-6}^1}=\frac{-27}{6}}\]

mathslover (mathslover):

got my point?

OpenStudy (anonymous):

4.5

mathslover (mathslover):

-4.5 ?

mathslover (mathslover):

check that there is - sign

OpenStudy (anonymous):

would that be x=-4.5?

mathslover (mathslover):

right ..

mathslover (mathslover):

now can u graph the following coordinates on the graph : (-4.5,0) and (0,-9)

OpenStudy (anonymous):

I think so, it would be x(-4.5,0) and y(0,-9) right?

mathslover (mathslover):

right

OpenStudy (anonymous):

ok Thanks so much....now one more question, how do I give you a medal?

mathslover (mathslover):

there is best response written?

OpenStudy (anonymous):

yes

mathslover (mathslover):

click on that like this

OpenStudy (anonymous):

ok done

mathslover (mathslover):

thanks

OpenStudy (anonymous):

Have a great day, I have to get back to work before I get fired lol (waving bye)